# Coloring and density theorems for configurations of a given volume

### v2

@@ 1.3. Parallelograms and parallelotopes
Unable to answer the question in Problem 6, we only state a partial result that will be obtained by modifying easily the ideas used for Theorem 3.
@@
As a related observation, Erdős remarked [25, p. 324] (without details or a reference) that he and Mauldin constructed a set $A\subseteq\mathbb{R}^{2}$ of infinite area that does not contain vertices of a parallelogram of area $1$ (also see the comments under Problem [6, #353]). Even if this has no implications to finite colorings of $\mathbb{R}^{2}$ or to positive density subsets of $\mathbb{R}^{2}$, it still hints that sets that avoid the vertices of parallelograms of unit area can be quite large. Without knowing which set Erdős and Mauldin had in mind, we observe that the example can be quite simple: an $(x,y)$-region between the positive parts of the coordinate axes $x=0$ and $y=0$ and the hyperbola $12xy=$ 1. In fact, we will easily prove the following stronger higher-dimensional result.
@@
Let $n\geqslant 2$ be a positive integer and take $\varepsilon>0$. There exists a Jordan-measurable set (i.e., a set with boundary of measure 00) $A\subseteq\mathbb{R}^{n}$ of infinite volume such that every $n$-dimensional parallelotope with all $2^{n}$ vertices in $A$ has volume less than $\varepsilon$.
@@ 1.4. Distance graphs in the plane
Here we discuss two lines of research that, at first, do not seem to be related to the paper topic, as they will be concerned with certain non-rigid structures, as opposed to polytopes. However, it will turn out they they are, in fact, related to Theorem 3 from Subsection 1.2 and to Graham’s theorem stated in Subsection 1.1, so they fit naturally in the present paper too. For this reason, unsurprisingly, we will be able to reuse the same techniques applied to our previous results, both positive and negative ones.
@@
Let $\mathcal{T}=(V,E)$ be a finite tree with the set of vertices $V$ and the set of edges $E$. To every edge $k\in E$ we assign a number $a_{k}\in(0,\infty)$ interpreted as edge-length, so we call the resulting structure a distance tree. It follows from a more general result by Lyall and Magyar [44, Theorem 2(i)] that
for every distance tree $\mathcal{T}$ and every set $A\subseteq\mathbb{R}^{2}$ of positive upper Banach density, there exists a number $\lambda_{0}>0$ (depending on $\mathcal{T}$ and $A$) such that for every number $\lambda\geqslant\lambda_{0}$ the set $A$ contains the vertices of an isometric copy of $\mathcal{T}$ scaled by the factor $\lambda$.
@@ 1.4.1. Embeddings of hypercube graphs
Embeddings of more general distance graphs are much more complicated and only partially resolved. Whenever the graph has a cycle, one can choose its edge-lengths to violate the triangle inequality, so it will not be embeddable in the Euclidean space. Lyall and Magyar [44] resolved this problem by starting from a distance graph that already is embedded in some other Euclidean space. Then they gave sufficient conditions for it to have all large dilates found in every set $A\subseteq\mathbb{R}^{d}$ of positive upper Banach density. However, even when their [44, Theorem 2] applies, it is rarely known what the minimal ambient dimension $d$ can be. Surprisingly, this is open already in the case of the equilateral (or any other non-degenerate) triangle.
@@
$[0,a_{1}]\times[0,a_{2}]\times\cdots\times[0,a_{n}].$ (1.2)
@@
This case of an embedding in $\mathbb{R}^{2}$ was not covered by [44, Theorem 2]; their result would require the ambient dimension to be at least $n+1$. Still, Predojević and the author [42, Theorem 1 and Remark 3] showed that
@@
There exists a Jordan-measurable finite coloring of $\mathbb{R}^{2}$ such that no color-class contains an embedding (1.3) of the $1$-skeleton of an $n$-box (1.2) satisfying $a_{1}\cdots a_{n}=1$.
In particular, a measurable set $A\subseteq\mathbb{R}^{2}$ with $\bar{\delta}_{2}(A)>0$ need not contain such an embedding. Theorem 8 will be shown by a simple modification of the construction used to prove Theorem 3 above. This might sound a bit surprising, since the quantity is $a_{1}\cdots a_{n}$ is the volume of the original box (1.2), but it has no such interpretation for an embedding of its $1$-skeleton (1.3).
@@ 1.4.2. Hyperbolic embeddings
Bardestani and Mallahi-Karai [5] studied a generalization of the Hadwiger–Nelson problem in which the points $v$ and $w$ of the same color are not allowed to satisfy the equation $Q(v-w)=1$, where $Q$ is a quadratic form over some local field. Under the additional measurability assumption, they were able to characterize when the corresponding chromatic number is infinite. Bardestani also informed the author that an easy modification of the argument from [5], which in turn uses a bound on the so-called independence ratio by Bachoc, DeCorte, de Oliveira Filho, and Vallentin [4], also implies a density variant of the same result in the plane $\mathbb{R}^{2}$ in the case of real hyperbolic forms; see Theorem 9 below. Davies [13] was motivated by a few unresolved questions in the particular hyperbolic case
@@
and its higher-dimensional space-time generalizations. He observed that Graham’s result on right triangles is relevant in that matter and it already implies that the corresponding (non-measurable) chromatic number is infinite as well.
It is natural to consider a coloring of a large square $[0,R]^{2}$ and impose bounds on $R$ sufficient to find equally colored points $(x,y)$ and $(x^{\prime},y^{\prime})$ such that
@@
    $\frac{|A|}{R^{2}}\geqslant\frac{1280}{\log R},$
@@
    If $R\in(0,\infty)$ and if the square $[0,R]^{2}$ is measurably colored in $r$ colors, then
    $R\geqslant e^{1280r}$
@@
    If $R\in(0,\infty)$ and if the square $[0,R]^{2}$ is measurably colored in $r$ colors, then
@@
    guarantees that there exist $n+1$ equally colored points of the form (1.5) for some vectors $v_{1},\ldots,v_{n}\in\mathscr{H}$.
@@
Part (b) is, once again, a clear consequence of part (a). We will use a similar outline and the same ideas employed in the proofs of Theorems 1 and 5 to establish Theorem 10 too. See Figure 3 for an illustration of the case $n=3$. Note that, in this hyperbolic setting, we are not allowing the freedom to scale the hyperbola $\mathscr{H}$ by some large factor $\lambda$. It is a non-compact curve, so it leaves more freedom than the circle to find non-scaled copies of the desired configuration. In the language of distance trees introduced at the beginning of Subsection 1.4, we are looking for certain hyperbolic embeddings of star trees (or simply stars).
@@ 1.5. Paper outline and the main ideas
Each of Sections 26 and 810 is dedicated to the proof of a single theorem. We first establish the “negative” results, since they are more elementary. Theorem 3 relies on an explicit partition of the complex plane $\mathbb{C}$, by coloring the number $z$ based on the relative location of $z^{2}$ to the (scaled) Gaussian integers $\mathbb{Z}+\mathbbm{i}\mathbb{Z}$. The construction could be thought of as a complex modification of the approach of Erdős at al. [27, §3] who used $|z|^{2}$ in place of $z^{2}$. The proof of Theorem 6 uses the same idea, only with roughly $1/\sin\varepsilon$ many colors, while the proof of Theorem 8 extends it to higher complex powers, namely $z^{n}$ for $n\geqslant 3$. The proof of Theorem 4 proceeds differently, by constructing the coloring that prevents monochromatic boxes that are only slightly rotated from their “standard” position and then using compactness of the group of rotations $\textup{SO}(n)$. The trick of using compactness of the underlying transformation group is essentially due to Straus [54], but its applicability is far from automatic: for instance, it can be used for rectangles but not for general parallelograms. Finally, Theorem 7 has a rather straightforward inductive proof, by estimating the determinant representing the volume of a parallelotope.
@@
Proofs of the “positive” density results rely on harmonic analysis, so a short Section 7 serves as an interlude, containing a few basic facts and some notation. The proof of Theorem 5 follows a similar outline as [20, §7] by Durcik and the author, which, in turn, stems from a general outline of performing a certain regularity decomposition of the counting form [7, 8, 11]. Similar proofs also appeared in a couple of other recent papers [40, 42], but here we also capitalize on the fact that the “fixed volume” condition is more flexible than fixing edge-lengths, so a lower ambient dimension suffices for that reason. The proof of Theorem 1 brings an additional novelty, by reinventing the notion of configuration size and defining it according to the lengths of some edges only, leaving the room for pigeonholing and achieving the unit volume by fixing the length of the remaining edge. In the proof of Theorem 10 we need to take care of several edges simultaneously, making it also similar to [40, §3].
@@
Section 11, the closing one, contains a short philosophical discussion on why area $1$ triangles allow a density theorem (in sufficiently large dimensions), while area $1$ rectangles do not. We draw analogies with singular integral forms from multilinear harmonic analysis.
@@
The author hopes that the present paper will revive the topic of studying existence theorems for fixed-volume point configurations, which seems to have been explored in only a few papers [47, 17, 3, 1, 48] since the early work of Graham [34].
@@ 2. Rectangles: proof of Theorem 3
In this definition we specify the vertex $A$ to be the one with the smallest coordinate $z_{A}$ in the lexicographic ordering of $\mathbb{C}\equiv\mathbb{R}^{2}$. Otherwise, $\mathscr{I}(\mathcal{P})$ would have only been determined up to multiplication by ±$\pm 1$.
@@
Consecutive side lengths of $\mathcal{P}$ are $|u|$ and $|v|$, so we have
@@
The above set does not intersect the circle (2.2); see Figure 5. Indeed, the central square lies fully inside (2.2) because of $24\sqrt{2}/3<$ 2, while all remaining open squares clearly belong to its exterior. ∎
@@ 3. Rectangular boxes: proof of Theorem 4
Here $\mathbb{S}_{m}$ denotes the group of permutations on the set $\{1,2,\ldots,m\}$. It is peculiar to notice that the right hand side RHS of (3.2) is the permanent of the matrix $(v_{j,k})_{j,k}$.
@@
$-p_{1}+(p_{1}+v_{1,1})=v_{1,1}.$
Take a positive integer $m\geqslant 2$. Observe that the left hand side LHS of (3.2) is a homogeneous polynomial of degree $m$ in $m^{2}+m$ variables $p_{k}$ and $v_{j,k}$, but the degree of each of the variables in it is at most $1$. Differentiating it with respect to the variable $v_{m,l}$ for some $l\in\{1,\ldots,m\}$ we obtain
@@
and renamed the indices. Therefore
$\textup{LHS}-\sum_{l=1}^{m}v_{m,l}\frac{\partial}{\partial v_{m,l}}\textup{LHS}=\textup{LHS}-\textup{RHS}$
and this quantity does not depend on $v_{m,1},\ldots,v_{m,m}$, so it can be evaluated simply by plugging $v_{m,1}=\cdots=v_{m,m}=0$ into the left hand side:
$\sum_{T\subseteq\{1,\ldots,m\}}(-1)^{m-|T|}\prod_{k=1}^{m}\Big{(}p_{k}+\sum_{\begin{subarray}{c}1\leqslant j\leqslant m-1\\ j\in T\end{subarray}}v_{j,k}\Big{)}=0.$
The last sum equals 00 because its terms can be paired to cancel each other: adding/subtracting the element $m$ to/from the set $T$ gives exactly the same term, only with the opposite signs. We have obtained $\textup{LHS}-\textup{RHS}=0$. ∎
@@
Note that only the first $m$ out of $n$ coordinates of a point ${x}$ appear in the product $x_{1}\cdots x_{m}$. Also, here $\text{par}({x})$ denotes the parity of a vertex ${x}$, which is computed as follows. Choose the base vertex of $\mathcal{R}$ to be the one with the smallest coordinate representation in the lexicographic ordering of $\mathbb{R}^{n}$. The parity of any vertex of $\mathcal{R}$ is defined to be the parity of its distance from the base vertex in the $1$-skeleton graph of $\mathcal{R}$. This number is either 00 or $1$ and it changes as we move from a vertex to its neighbor along a $1$-edge of $\mathcal{R}$. Clearly, the expression (3.3) has $2^{m}$ terms, half of them get the $+$ sign and half of them come with the -- sign; see the illustration of these signs in Figure 7.
Figure 7. Signs attached to box vertices.
@@
where ${p}=U{q}$ and ${v}_{j}=a_{j}U\mathbbm{e}_{j}$ for each index $1\leqslant j\leqslant m$. Also, write coordinate-wise:
@@
for $10\leqslant l\leqslant 3\cdot 2^{m}-$ 1. We claim that the vertices of $\mathcal{R}$ cannot all belong to the same set $\mathscr{S}_{l}$. Namely, if they did, then the definition of $\mathscr{J}(\mathcal{R})$ would give
@@
where $(l_{1},l_{2},\ldots,l_{L})$ run over all $L$-tuples of elements from $\{0,1,2,\ldots,3\cdot 2^{m}-1\}$. In words, this is a joint refinement of $L$ mutually rotated colorings. Suppose that the vertices of $\mathcal{R}$ belong to the same color-class $\mathscr{C}_{l_{1},l_{2},\ldots,l_{L}}$. Let $U\in\textup{SO}(n)$ be as in (3.4), but without any assumption on the norm of $U-I$. Take an index $i\in\{1,\ldots,L\}$ such that $U\in U_{i}\mathcal{O}$. Then the box
@@
The number of colors needed in the above proof grows superexponentially in $n$.

### v3

@@ 1.3. Parallelograms and parallelotopes
Unable to answer the question in Problem 6, we only state a partial result that will be obtained by easily modifying the ideas used for Theorem 3.
@@
As a related observation, Erdős remarked [26, p. 324] (without details or a reference) that he and Mauldin constructed a set $A\subseteq\mathbb{R}^{2}$ of infinite area that does not contain vertices of a parallelogram of area $1$ (also see the comments under Problem [6, #353]). Even if this has no implications to finite colorings of $\mathbb{R}^{2}$ or to positive density subsets of $\mathbb{R}^{2}$, it still hints that sets that avoid the vertices of parallelograms of unit area can be quite large. Without knowing which set Erdős and Mauldin had in mind, we observe that the example can be quite simple: an $(x,y)$-region between the positive parts of the coordinate axes $x=0$ and $y=0$ and the hyperbola $2xy=1$. In fact, we will easily prove the following stronger higher-dimensional result.
@@
Let $n\geqslant 2$ be a positive integer and take $\varepsilon>0$. There exists a Jordan-measurable set (i.e., a set with boundary of measure 0) $A\subseteq\mathbb{R}^{n}$ of infinite volume such that every $n$-dimensional parallelotope with all $2^{n}$ vertices in $A$ has volume less than $\varepsilon$.
@@ 1.4. Distance graphs in the plane
Here we discuss two lines of research that, at first, do not seem to be related to the paper topic, as they will be concerned with certain non-rigid configurations, as opposed to polytopes. However, it will turn out, quite surprisingly, that Theorem 8 below can be proved simply by modifying the proof of Theorem 3 from Subsection 1.2, while the configuration of interest in Theorem 9 below will be obtained by taking two out of three vertices of Graham’s axes-aligned triangle from Subsection 1.1 and performing a clever change of variables appearing in [5] and [14]. Thus, the study of these flexible configurations fits naturally to the present paper as well and we will be able to reuse the same techniques applied to our previous results, both positive and negative ones.
@@
Let $\mathcal{T}=(V,E)$ be a finite tree with the set of vertices $V$ and the set of edges $E$. To every edge $k\in E$ we assign a number $a_{k}\in(0,\infty)$ interpreted as edge-length, so we call the resulting structure a distance tree. It follows from a more general result by Lyall and Magyar [48, Theorem 2(i)] that
for every distance tree $\mathcal{T}$ and every set $A\subseteq\mathbb{R}^{2}$ of positive upper Banach density, there exists a number $\lambda_{0}>0$ (depending on $\mathcal{T}$ and $A$) such that for every number $\lambda\geqslant\lambda_{0}$ the set $A$ contains the vertices of an isometric copy of $\mathcal{T}$ scaled by factor $\lambda$.
@@ 1.4.1. Embeddings of hypercube graphs
Embeddings of more general distance graphs are much more complicated and only partially resolved. Whenever the graph has a cycle, one can choose its edge-lengths to violate the triangle inequality, so it will not be embeddable in the Euclidean space. Lyall and Magyar [48] resolved this problem by starting from a distance graph that already is embedded in some other Euclidean space. Then they gave sufficient conditions for it to have all large dilates found in every set $A\subseteq\mathbb{R}^{d}$ of positive upper Banach density. However, even when their [48, Theorem 2] applies, it is rarely known what the minimal ambient dimension $d$ can be. Surprisingly, this is open already in the case of the equilateral (or any other non-degenerate) triangle.
@@
$0,a_{1}]\times[0,a_{2}]\times\cdots\times[0,a_{n}].$ (1.2)
@@
This case of an embedding in $\mathbb{R}^{2}$ was not covered by [48, Theorem 2]; their result would require the ambient dimension to be at least $n+1$. Still, Predojević and the author [45, Theorem 1 and Remark 3] showed that
@@
Fix a positive integer $n$. There exists a Jordan-measurable finite coloring of $\mathbb{R}^{2}$ such that no color-class contains an embedding (1.3) of the $1$-skeleton of an $n$-box (1.2) satisfying $a_{1}\cdots a_{n}=1$.
In particular, a measurable set $A\subseteq\mathbb{R}^{2}$ with $\bar{\delta}_{2}(A)>0$ need not contain such an embedding. Theorem 8 will be shown by a simple modification of the construction used to prove Theorem 3 above. This might sound a bit surprising, since the quantity $a_{1}\cdots a_{n}$ is the volume of the original box (1.2), but it has no such interpretation for an embedding of its $1$-skeleton (1.3). This, however, leads to structural differences between embeddings of parallelotopes of unit volume and their partial skeletons with edge-lengths multiplying to $1$, but we provided negative results for both. Unit volume rectangular boxes lie at their intersection, which enabled us to obtain the strongest result on their avoidance, namely Theorem 4.
@@ 1.4.2. Hyperbolic embeddings
Bardestani and Mallahi-Karai [5] studied a generalization of the Hadwiger–Nelson problem in which the points $v$ and $w$ of the same color are not allowed to satisfy the equation Q(vw)=1Q(v-w)=1, where $Q$ is a quadratic form over some local field. Under the additional measurability assumption, they were able to characterize when the corresponding chromatic number is infinite. Bardestani also informed the author that an easy modification of the argument from [5], which in turn uses a bound on the so-called independence ratio by Bachoc, DeCorte, de Oliveira Filho, and Vallentin [4], also implies a density variant of the same result in the plane $\mathbb{R}^{2}$ in the case of real hyperbolic forms; see Theorem 9 below. Davies [14] was motivated by a few unresolved questions in the particular hyperbolic case
@@
and its higher-dimensional space-time generalizations. He observed that Graham’s result on right triangles is relevant in that matter and it already implies that the corresponding (non-measurable) chromatic number is infinite as well. Namely, if $(u,v)$, $(u^{\prime},v^{\prime})$, $(u^{\prime},v)$ are vertices of an area $1$ right triangle with $u>u^{\prime}$ and $v<v^{\prime}$, then the points
$\begin{pmatrix}x\\ y\end{pmatrix}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}\begin{pmatrix}u\\ v\end{pmatrix},\quad\begin{pmatrix}x^{\prime}\\ y^{\prime}\end{pmatrix}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}\begin{pmatrix}u^{\prime}\\ v^{\prime}\end{pmatrix}$
satisfy $Q((x,y),(x^{\prime},y^{\prime}))=1$.
It is natural to consider a coloring of a large square $0,R]^{2}$ and impose bounds on $R$ sufficient to find equally colored points $(x,y)$ and $(x^{\prime},y^{\prime})$ such that
@@
There exists a constant $C\in(0,\infty)$ with the following properties.
@@
    $\frac{|A|}{R^{2}}\geqslant\frac{C}{\log R},$
@@
    If $R\in(0,\infty)$ and if the square $0,R]^{2}$ is measurably colored in $r$ colors, then
    $R\geqslant e^{Cr}$
@@
    If $R\in(0,\infty)$ and if the square $0,R]^{2}$ is measurably colored in $r$ colors, then
@@
    guarantees that there exist $n+1$ equally colored points of the form (1.5) for some $z\in\mathbb{R}^{2}$ and vectors $v_{1},\ldots,v_{n}\in\mathscr{H}$.
@@
In the language of distance trees introduced at the beginning of Subsection 1.4, we are looking for certain hyperbolic embeddings of star trees (or simply stars). Part (b) is, once again, a clear consequence of part (a). We will use a similar outline and the same ideas employed in the proofs of Theorems 1 and 5 to establish Theorem 10 too. See Figure 3 for an illustration of the case $n=3$.
Note that, in this hyperbolic setting, we are not allowing the scaling of the hyperbola $\mathscr{H}$ by some factor $\lambda$. It is a non-compact curve, so it leaves more freedom than the circle to find non-scaled copies of the desired configuration. For this reason, as it was discussed in [5], the result analogous to Theorem 10 is possible for every indefinite quadratic form in place of $Q(x,y)=x^{2}-y^{2}$, but not for semi-definite ones.
Remark 11.
@@
To avoid possible confusion, let us emphasize that this subsection did not study configurations determined by fixing some hyperbolic distances—it was “hyperbolic” rather in the sense of the Minkowski spacetime. Adapting our techniques to hyperbolic geometry would be a different task that we do not attempt in the present paper.
@@ 1.5. Paper outline and the main ideas
Each of Sections 26 and 810 is dedicated to the proof of a single theorem. We first establish the “negative” results, since they are more elementary. Theorem 3 relies on an explicit partition of the complex plane $\mathbb{C}$, coloring the number $z$ based on the relative location of $z^{2}$ to the (scaled) Gaussian integers $\mathbb{Z}+\mathbbm{i}\mathbb{Z}$. The construction could be thought of as a complex modification of the approach of Erdős at al. [28, §3] who used $|z|^{2}$ in place of $z^{2}$. The proof of Theorem 6 uses the same idea, only with roughly $1/\sin\varepsilon$ many colors, while the proof of Theorem 8 extends it to higher complex powers, namely $z^{n}$ for $n\geqslant 3$. The proof of Theorem 4 proceeds differently, by constructing the coloring that prevents monochromatic boxes that are only slightly rotated from their “standard” position and then using compactness of the group of rotations $\textup{SO}(n)$. One could say that the main novelty here is the construction of an almost invariant quantity for the boxes of the aforementioned type and not insisting on finding the exact invariant, which might not even exist in full generality. The trick of using compactness of the underlying transformation group is essentially due to Straus [59], but in a somewhat different setting. Also, its applicability is far from automatic: for instance, it can be used for rectangles but not for general parallelograms. Finally, Theorem 7 has a rather straightforward inductive proof, by estimating the determinant representing the volume of a parallelotope.
@@
Proofs of the “positive” density results rely on harmonic analysis, so a short Section 7 serves as an interlude, containing a few basic facts and some notation. The proof of Theorem 5 follows a similar outline as [20, §7] by Durcik and the author, which, in turn, stems from a general outline of performing a certain regularity decomposition of the counting form [7, 8, 11]. Similar proofs also appeared in a couple of other recent papers [43, 45], but here we also capitalize on the fact that the “fixed volume” condition is more flexible than fixing edge-lengths, so a lower ambient dimension suffices for that reason. The proof of Theorem 1 brings an additional novelty, by reinventing the notion of configuration size and defining it according to the lengths of some edges only, leaving room for pigeonholing and achieving the unit volume by fixing the length of the remaining edge. In the proof of Theorem 10 we need to take care of several edges simultaneously, making it also similar to [43, §3]. Here we are making use of the fact that the underlying configuration is no longer rigid but flexible, which gives us additional useful “degrees of freedom” to detect the configuration analytically.
@@
Section 11 contains a short philosophical discussion on why area $1$ triangles allow a density theorem (in sufficiently large dimensions), while area $1$ rectangles do not. We draw analogies with singular integral forms from multilinear harmonic analysis. Section 12, the closing one, contains a simplified table summarizing both known and open results for several basic types of point configurations.
@@
The author hopes that the present paper will revive the topic of studying existence theorems for fixed-volume point configurations, which seems to have been explored in only a few papers like [51, 17, 3, 1, 52, 46, 40] since the early work of Graham [36].
@@ 2. Rectangles: proof of Theorem 3
In this definition we specify the vertex $A$ to be the one with the smallest coordinate $z_{A}$ in the lexicographic ordering of $\mathbb{C}\equiv\mathbb{R}^{2}$. Otherwise, $\mathscr{I}(\mathcal{P})$ would have only been determined up to multiplication by $\pm 1$.
@@
Consecutive side-lengths of $\mathcal{P}$ are $|u|$ and $|v|$, so we have
@@
The above set does not intersect the circle (2.2); see Figure 5. Indeed, the central square lies fully inside (2.2) because of $4\sqrt{2}/3<2$, while all remaining open squares clearly belong to its exterior. ∎
@@
The above proof of Theorem 3 has recently been formalized in Lean [6, #189] with the help of Aristotle AI (by Harmonic) and Gemini 3 Pro (by Google), as part of joint efforts of the Erdős problems community [6].
@@ 3. Rectangular boxes: proof of Theorem 4
Here $\mathbb{S}_{m}$ denotes the group of permutations on the set $\{1,2,\ldots,m\}$. It is peculiar to notice that the right hand side (RHS) of (3.2) is the permanent of the matrix $(v_{j,k})_{j,k}$. In fact, identity (3.2) is a generalization of Ryser’s formula for the permanent [53, Corollary 4.2], which is obtained by specializing $p_{k}=0$ for every $k$. We will give a different self-contained proof.
@@
$p_{1}+(p_{1}+v_{1,1})=v_{1,1}.$
Take a positive integer $m\geqslant 2$. Observe that the left hand side (LHS) of (3.2) is a homogeneous polynomial of degree $m$ in $m^{2}+m$ variables $p_{k}$ and $v_{j,k}$, but the degree of each of the variables in it is at most $1$. Differentiating it with respect to the variable $v_{m,l}$ for some $l\in\{1,\ldots,m\}$ we obtain
@@
and renamed the indices. Multiplying by $v_{m,l}$ and summing in $l$ gives
$\sum_{l=1}^{m}v_{m,l}\frac{\partial}{\partial v_{m,l}}\textup{LHS}=\textup{RHS}.$
Observe that the polynomial
$\textup{LHS}-\sum_{l=1}^{m}v_{m,l}\frac{\partial}{\partial v_{m,l}}\textup{LHS}$
does not depend on the variables $v_{m,1},\ldots,v_{m,m}$, because they appear with power at most $1$ and thus certainly cancel when evaluating the above difference. Therefore, the quantity $\textup{LHS}-\textup{RHS}$ also does not depend on $v_{m,1},\ldots,v_{m,m}$, so it can be evaluated simply by plugging $v_{m,1}=\cdots=v_{m,m}=0$:
$\textup{LHS}-\textup{RHS}=\sum_{T\subseteq\{1,\ldots,m\}}(-1)^{m-|T|}\prod_{k=1}^{m}\Big(p_{k}+\sum_{\begin{subarray}{c}1\leqslant j\leqslant m-1\\ j\in T\end{subarray}}v_{j,k}\Big)=0.$
The last sum equals 0 because its terms can be paired to cancel each other: adding/subtracting the element $m$ to/from the set $T$ gives exactly the same term, only with the opposite signs. We have obtained $\textup{LHS}=\textup{RHS}$. ∎
@@
Note that only the first $m$ out of $n$ coordinates of a point x{x} appear in the product $x_{1}\cdots x_{m}$. Also, here $\text{par}({x})$ denotes the parity of a vertex x{x}, which is computed as follows. Choose the base vertex of $\mathcal{R}$ to be the one with the smallest coordinate representation in the lexicographic ordering of $\mathbb{R}^{n}$. The parity of any vertex of $\mathcal{R}$ is defined to be the parity of its distance from the base vertex in the $1$-skeleton graph of $\mathcal{R}$. This number is either 0 or $1$ and it changes as we move from a vertex to its neighbor along a $1$-edge of $\mathcal{R}$. Clearly, the expression (3.3) has $2^{m}$ terms, half of them get the ++ sign and half of them come with the - sign; see the illustration of these signs in Figure 7.
Figure 7. Signs attached to rectangle and box vertices (cases $m=2$ and $m=3$).
@@
where p=Uq{p}=U{q} and ${v}_{j}=a_{j}U\mathbbm{e}_{j}$ for each index $1\leqslant j\leqslant m$. Also, write coordinatewise:
@@
for $0\leqslant l\leqslant 3\cdot 2^{m}-1$. We claim that the vertices of $\mathcal{R}$ cannot all belong to the same set $\mathscr{S}_{l}$. Namely, if they did, then the definition of $\mathscr{J}(\mathcal{R})$ would give
@@
where $(l_{1},l_{2},\ldots,l_{L})$ run over all $L$-tuples of elements from $\{0,1,2,\ldots,3\cdot 2^{m}-1\}$. In words, this is a joint refinement of $L$ mutually rotated colorings. Suppose that the vertices of $\mathcal{R}$ belong to the same color-class $\mathscr{C}_{l_{1},l_{2},\ldots,l_{L}}$. Let $U\in\textup{SO}(n)$ be as in (3.4), but without any assumption on the norm of UIU-I. Take an index $i\in\{1,\ldots,L\}$ such that $U\in U_{i}\mathcal{O}$. Then the box
@@
The number of colors needed in the above proof grows superexponentially in $n$, as remarked in the introduction. This follows from the known growth of the minimal cardinality $L(n)$ of a cover of $\textup{SO}(n)$ by translates of the set $\mathcal{O}$ (or of any other fixed small neighborhood of $I$). Namely, when the Haar measure $\mu_{\textup{SO}(n)}$ of the orthogonal group is normalized to have total mass $1$, then the large deviation principle for the spectral measure of the orthogonal group estimates $\mu_{\textup{SO}(n)}(\mathcal{O})=O((\varepsilon/2)^{n^{2}/2})$ as $n\to\infty$ for a fixed $\varepsilon\in(0,1)$; we omit the lengthy details. Therefore, it remains to apply the “volume bound”
$L(n)\geqslant\frac{\mu_{\textup{SO}(n)}(\textup{SO}(n))}{\mu_{\textup{SO}(n)}(\mathcal{O})}$
to conclude the superexponential growth of $L(n)$ in $n$.